Obavijest

U okviru Seminara za matematičku logiku i osnove matematike, u \underline{ponedjeljak, 23. svibnja 2005.}, mr. Ružica Piskač, University of Saarbruecken održat će predavanje pod naslovom

The Verification of a Checker for Priority Queues.

Predavanje će se održati u predavaonici 203 PMF-MO, s početkom u \underline{17:15} sati.

\underline{Sažetak:}

We formally prove the correctness of the time superefficient result checker for priority queues, which is implemented in LEDA. A priority queue is a data structure that supports insertion, deletion and retrieval of the minimal element, relative to some order.

A result checker for priority queues is a data structure that monitors the input and output of the priority queue. Whenever the user requests a minimal element, it checks that the returned element is indeed minimal. In order to do this, the checker makes use of a system of lower bounds.

We have verified that, for every execution sequence in which the checker accepts the outputs, the priority queue returned the correct minimal elements. For the formal verification, we used the firstorder theorem prover Saturate.

Tajnik seminara: Vedran Čačić